A boy is standing atop a staircase of 8 stairs each stair being 20 cm high and 20 cm wide. He releases a ball of 500 gm from 50 cm above the floor. 1) what is kinetic energy at the time of release of ball? 2) what will be velocity of ball at third step from ground? (g= 10 m/s^2)
1
Expert's answer
2021-10-21T10:53:25-0400

1)

"v=0,"

"E=\frac{mv^2}2=0~\text J,"

2)

"v=\sqrt{2gh}=\sqrt{2\cdot 10\cdot \frac 38\cdot 0.5}=3.75~\frac ms."